OpenStudy (canimcan):

Hello to everyone. :) I need help with this... Determine the equation of surface which is formed by the tangents of ellipsoid x^2 + y^2/2+z^2=1. Also that surface contain point M(5,5,5).

OpenStudy (canimcan):

Actually I need to get this equation 123[x^2 + y^2/2 +z^2 - 1]=50(x+ y/2 + z- 1/5]^2 Thank you for suggestions!

OpenStudy (3mar):

Does your equation take that form? \[123(x^2+\frac{ y^2 }{ 2 }+z^2-1)=50(x+\frac{ y }{ 2 }+z-\frac{ 1 }{ 5 })^2\]

OpenStudy (canimcan):

Yes.That is result.How to get that form?

OpenStudy (3mar):

Click on Equation button under the text area

OpenStudy (canimcan):

how to solve this problem I mean.....

OpenStudy (3mar):

If you want to solve such equation, it is not possible as you have three variables, so you must have three equations to get their values. n- variables ....... n- equations.

OpenStudy (3mar):

But if you want to konw how to type such form, it is possible and easy. Just click on the button "Equation" and choose what you want to express. Try more and more times, you will master it.